OpenStudy (anonymous):

2x +3y = 2 x+6y = 4 Solve

OpenStudy (anonymous):

re-arrange the second equation (most easiest equation) in terms of x: x=4-6y sub this into the first equation 2x+3y=2 2(4-6y)+3y=2 expand this and solve for y, then use this y value and plug it back into x=4-6y

OpenStudy (anonymous):

2x+3y=2---(1) x+6y=4----(2) from (1) equation we get 2x=2-3y =>x=2-3y/2---(3) now we should replace the value of x in number (2) equation so,2-3y/2+6y=4 =>2-3y+12y/2=4 =>2-3y+12y=8 =>9y=8-2 =>9y=6 =>y=6/9 =>y=2/3 by using value of y from (3) equation we get?

OpenStudy (anonymous):

multiply first equation with 2 then minus them ie 4X+6Y= 4 X+6Y= 4 so we get (minus it) 3X =0 X =0 Y = 4/6 =2/3

OpenStudy (luigi0210):

@Velmina are you even trying or are you just waiting for the answer?

OpenStudy (anonymous):

I've been trying

OpenStudy (anonymous):

If I could do it I would not be on here asking